OpenStudy (anonymous):

a pound of of u.s mud contains an average of 7.55X10^11 bacteria. there are 60.0 drops in one teaspoon, 6.0 teaspoons in an ounce, and 16 oz in a pound. how many bacteria live in a drop of mud?

OpenStudy (turingtest):

\[\frac{7.55\times10^{11}\text{bact}}{\text{lb}}\times\frac{\text{lb}}{16\text{oz}}\times\frac{\text{oz}}{6\text{tsp}}\times\frac{\text{tsp}}{60\text{drops}}=?\frac{\text{bacteria}}{\text{drop}}\]

OpenStudy (anonymous):

ok ty